# SALAD: Achieve High-Sparsity Attention via Efficient Linear Attention Tuning for Video Diffusion Transformer

### v1

@@ A.2.1 Training Settings
Table 6: Implementation of Salad Tuning
optimizer AdamW
learning rate 1e-4
training steps 1600
batch size 8
number of training samples 2000
Total tuning GPU hours 20.6
A.2.2 Training Cost
SALAD requires only 2,000 training videos and approximately 20.6 GPU hours with a batch size of 8 and achieves effective high-sparsity adaptation in Video Diffusion Transformers. This constitutes a substantially lower computational footprint compared with methods that pretrain sparse-attention architectures from scratch.
For reference, VSA [37] is pretrained on 80,000 videos using 32 NVIDIA H100 GPUs, while VMoBA [23] leverages the Koala-36M dataset (36 million video clips) and 104 NVIDIA H800 GPUs under the token sequence length of 33K. SLA [34], a recent linear-attention-based approach for high-sparsity adaptation, is tuned on 20,000 self-collected videos—ten times more than SALAD. SLA also requires a batch size of 64 for tuning. Under our experimental setting, SLA fails to recover the performance of the original dense model as reported in Tab. 1, whereas SALAD achieves comparable or superior results at a fraction of the resource cost. These findings underscore SALAD’s tuning efficiency for sparse adaptation in large-scale video diffusion models.
Method Subject Background Image Text Trainable
Consistency Consistency Quality Consistency Parameters
Original (Full Attention) 94.19 96.76 70.17 28.95 -
Spatial-Temporal Sliding Window Attention
LoRA (r = 64) 96.89 96.18 65.65 27.60 47M
LoRA (r = 128) 96.77 96.73 67.80 26.87 94M
LoRA (r = 256) 90.54 96.25 63.75 25.25 189M
LoRA (r = 512) 91.30 94.31 66.39 26.19 377M
SALAD 97.21 96.83 69.41 25.56 165M
TopK Sparse Attention (k=4)
LoRA (r = 128) 95.37 96.72 67.30 27.33 94M
LoRA (r = 256) 98.35 96.88 69.99 26.88 189M
LoRA (r = 512) 96.52 96.52 69.56 24.27 377M
SALAD 98.16 95.93 70.33 27.61 165M
A.3 Details of Post-training Branch Dropping
In Sec. 5.4 we mention that some linear branches can be dropped while maintaining performance. We now provide more details about it.
We first analyze the distribution of gate values across denoising timesteps. As illustrated in Fig. 9, the gate value distributions remain consistent across layers and timesteps. In particular, the 20th, 40th, 60th, and 80th percentiles vary smoothly over time. Since the gate value reflects, to some extent, the relative importance of each branch, we can set a threshold to determine when the linear-attention branch can be safely dropped. This observation motivates a simple yet effective thresholding strategy that uses the average percentile as the threshold, avoiding the need for per-timestep calibration.
Figure 9: Quintile Trends of Gate Values across Time-steps. The distributions of gate values remain remarkably stable throughout the diffusion process. This temporal consistency motivates our timestep-agnostic thresholding strategy for branch selection.
Using the above thresholds, we conducted a series of experiments to investigate the impact of dropping branches according to their gate values. Specifically, we progressively drop linear attention branches in ascending order of gate magnitude. Fig. 10 reports the performance results on VBench subset. Notably, dropping approximately 20% of the branches led to the optimal video generation metrics. Increasing the proportion of dropped branches beyond this point caused a decline in quality, indicating that excessive pruning compromises the model’s expressive capacity.
Figure 10: Progressive Branch Dropping Analysis. IQ (Image Quality) scores across different branch reservation strategies. The shaded region shows the effective dropping range that maintains baseline quality with around 20% branch dropped.
To further figure out the best dropping plan, we performed an ablation study by dropping branches from different gate intervals. Tab. 8 summarizes these results. Removing the branches corresponding to the largest 20% of gate values consistently yielded superior metrics, especially for subject consistency, compared to dropping branches from other intervals or random selection, which may be partially attributed to the low-rank nature of these high-gate branches. By selectively dropping these branches, we not only preserve video quality but also achieve additional computational acceleration.
Table 8: Qualitative results with different drop strategies
Drop Strategy SC (↑) BC (↑) IQ (↑) TC (↑)
Original 97.21 96.83 69.41 25.56
0-20% 96.88 95.60 69.37 25.14
20%-40% 96.93 96.30 67.11 23.56
40%-60% 96.75 96.56 67.97 24.33
60%-80% 96.84 96.94 68.25 24.90
[…]
Random (20%) 96.89 96.27 66.79 24.40
A.4 Details of Non-Shared Sparse-Linear Attention
@@
For the experiments in Tab. 15, we follow the same training configuration for both shared and non-shared variants and evaluate all models on the same VBench subset. Though has more weights, the performance of naive non-shared sparse-linear attention does not significantly increase when compared with the shared sparse-linear attention variant. After incorporating the input-dependent gate and the zero-initialization strategy, its performance on all the metrics improves substantially.
@@
A.5 Observation on Input-Dependent Scalar Gate
Figure 11: Gate design. The computation of GATE.
Figure 12: Comparison of Branch and Sparse Attention Outputs. The top panel demonstrates the value range of the Sparse Attention Outputs and Linear Attention Branch Outputs. The blue shaded region denotes the sparse attention outputs, while the red shaded region highlights the gated branch outputs, which provide fine-grained complementary information without disrupting the original structural patterns. The bottom panel further visualizes the gate values, typically ranging between 0.1 and 0.4, indicating how the gating mechanism adaptively modulates the contribution of linear attention across different blocks.
As shown in Sec. A.3, the learned gate values range from 0 to 1, which constrains the contribution of the linear branch to a fine-grained interval, thereby balancing the overall distribution. As illustrated in Fig. 12, the magnitudes of the original Sparse Attention Output and the Linear Attention Branch Output differ substantially while the gate provides fine-grained modulation, complementing the information from the sparse attention while preserving its original structure.
To further investigate the role of the gate during inference, we conducted a controlled study by fixing its value between 0 and 1.5 on a model trained with a gate of 1. As shown in Fig. 13, varying the gate reveals distinct failure modes. Setting gate=0 forces the model to rely solely on the LoRA branch, resulting in severe color distortions (e.g., the dog’s fur appears unnaturally green or red), indicating that LoRA alone cannot recover global interactions. Increasing the gate to 0.7 partially activates linear attention, which improves overall sharpness but introduces semantic errors, such as duplicating the dog (e.g., two dogs appear in the scene instead of one). At gate=1, the output aligns with the prompt semantics, yet local spatial inconsistencies persist, such as mismatched head–body placement of the dog (e.g., disproportionate size of the dog’s head and body). Pushing the gate to 1.5 over-amplifies linear attention, causing generation collapse, producing blurred frames with noise. These observations suggest that the gate must be carefully moderated. A balanced linear attention contribution enriches sparse attention with complementary global cues while preserving the structural integrity of the generated content. By applying an input-dependent gate, such a balance of linear attention and sparse attention can be automatically learned from training.
Figure 13: Video example of the effect of Gate choices when inference. The model is trained with a gate value of 1 and evaluated under inference with different gate settings. Using gate=0 leads to severe color distortion, while gate=0.7 incorrectly introduces an additional dog. Although gate=1 aligns with the prompt semantics, it produces spatially inconsistent dog head–body structures. With gate=1.5, the model fails to generate meaningful outputs. In contrast, input-dependent gate values yield high-quality and semantically aligned results.
Figure 14: Effect of Gate Choices During Training. We compare models trained with fixed gate values to those trained with input-dependent gates. Fixed-gate models exhibit distinct semantic errors, such as gate=0 producing two dogs and gate=0.5 generating duplicate tails. Even gate=1 shows spatial inconsistencies. In contrast, the input-dependent gate model produces natural videos with coherent structures and fine-grained details (e.g., wagging tail).
A similar pattern arises during training. As illustrated in Fig. 14, constraining the gate to a fixed value throughout optimization substantially impairs model quality. With gate=0 (LoRA only), the model fails to capture semantic information and spatial correlation, as generates 2 dogs instead of 1. Setting gate=0.5 introduces semantic errors too, such as duplicated tails or misaligned subjects, indicating that fixed gates cannot adapt to varying contextual complexity across training samples. In contrast, the input-dependent gate dynamically modulates the contribution of the linear branch relative to sparse attention, enhancing global coherence while preserving fine-grained structural details (e.g., smoothly moving tails and consistent object placement). This content-aware, adaptive modulation is crucial for achieving both visual fidelity and stable training behavior.
These observations underscore the necessity of an input-dependent, dynamic gating mechanism to balance contributions from the sparse and linear branches, ensuring both structural fidelity and fine-grained detail preservation in both training and inference.
A.6 Ablation on the Non-Linear Function in the GATE
In this section, we investigate the choice of the non-linear function used in the GATE. As illustrated in the main paper and the Fig. 11, we introduce a non-linearity to increase the expressive capacity of the GATE, and we adopt the sigmoid function since a gating range constrained to $[0,1]$ empirically improves the performance of SALAD. Tab. 10 reports the ablation results across different non-linear functions. Sigmoid yields the best performance, while ReLU and Tanh underperform. This suggests that restricting the gate values to $[0,1]$ is sufficient and beneficial, which is also consistent with our visualization results.
Table 10: Comparison of Non-Linear Functions in the GATE
Non-Linear Function SC (↑) BC (↑) IQ (↑) TC (↑)
Tanh 96.89 96.25 66.20 26.97
ReLU 96.77 96.10 67.12 28.26
[…]
A.7 More Qualitative Results
@@
To further assess the visual quality and temporal consistency achieved by SALAD, we present additional qualitative comparisons against several representative baselines in this section. We evaluate SALAD with other baselines on Wan 2.1-1.3B. As illustrated in Figure 16 and Figure 17, SALAD consistently preserves high-fidelity appearance details while generating temporally coherent motion, often matching or surpassing the visual quality produced by dense-attention counterparts. Additional video examples are provided in the supplementary materials for a more comprehensive evaluation.

### v2

@@ 0.A.4 Ablation on the Non-Linear Function in the Input-Timestep-Aware Scaling Module
0.A.5 Details of Non-Shared Sparse-Linear Attention
@@
For the experiments in Tab. 13, we follow the same training configuration for both shared and non-shared variants and evaluate all models on the same VBench subset. Though has more weights, the performance of naive non-shared sparse-linear attention does not significantly increase when compared with the shared sparse-linear attention variant. After incorporating the input-dependent Dynamic Scaler and the zero-initialization strategy, its performance on all the metrics improves substantially.
@@
0.A.6 Extended Comparison with LoRA
In addition to the comparison presented in the main paper, we provide a more detailed evaluation of LoRA under different sparsity strategies in Tab. 9. Specifically, we report additional results for LoRA with multiple ranks under the ST-SWA and TopK ($k=4$) sparsity strategy.
Consistent with the observations in the main text, increasing the LoRA rank does not consistently improve generation quality. The best-performing LoRA configuration generally occurs at moderate ranks (e.g., $r=128$), while larger ranks can even degrade performance. In contrast, SALAD achieves competitive or better performance across most metrics, demonstrating that the improvements primarily arise from the design of the linear attention branch rather than simply increasing the number of tunable parameters.
Note that the trainable parameters reported in Tab. 9 include the parameters introduced by LoRA and linear attention branch during training. However, LoRA weights are merged into the base model after training and therefore do not introduce additional parameters during inference. Therefore, the additional parameter overhead reported in the contributions (4.99%) refers to the persistent parameters required by SALAD during inference.
Method Subject Background Image Text Trainable
Consistency Consistency Quality Consistency Parameters
Original (Full Attention) 94.19 96.76 70.17 28.95 -
ST-SWA
LoRA (r = 64) 96.89 96.18 65.65 27.60 47M
LoRA (r = 128) 96.77 96.73 67.80 26.87 94M
LoRA (r = 256) 90.54 96.25 63.75 25.25 189M
LoRA (r = 512) 91.30 94.31 66.39 26.19 377M
SALAD 97.21 96.83 69.41 25.56 165M
TopK (k=4)
LoRA (r = 128) 95.37 96.72 67.30 27.33 94M
LoRA (r = 256) 98.35 96.88 69.99 26.88 189M
LoRA (r = 512) 96.52 96.52 69.56 24.27 377M
SALAD 98.16 95.93 70.33 27.61 165M
0.A.7 Training Details
0.A.7.1 Training Settings
All training configurations are summarized in Tab. 10. We conduct experiments on the Wan2.1 family of video diffusion transformers [wan2025]. Unless otherwise specified, experiments are performed on Wan2.1-1.3B at a resolution of $480\times 832$ with 77 frames (approximately 30k tokens). We also extend Wan2.1-1.3B to a higher resolution setting of 720p (about 100k tokens) and apply SALAD to Wan2.1-14B at 480p.
All training-based methods are fine-tuned using 2,000 videos sampled from the Mixkit dataset under the OpenSora training protocol [lin2024open]. We apply LoRA to the attention projection layers ($W^{Q}$, $W^{K}$, $W^{V}$, $W^{O}$) with rank $r=128$ and $\alpha=256$. For SALAD, we additionally fine-tune the scaling-related modules (see Fig. 3).
Training is performed on 4 GPUs with a total batch size of 8 and a learning rate of $1\mathrm{e}{-4}$. For Wan2.1-1.3B at 480p, training runs for 1,600 steps. For the higher-resolution 720p setting, we train for 1,000 steps. For Wan2.1-14B at 480p, training requires 600 steps.
Table 10: Training details for SALAD fine-tuning under different settings. All experiments use 2,000 videos from the Mixkit dataset under the OpenSora training protocol [lin2024open], LoRA on attention projection layers ($W^{Q}$, $W^{K}$, $W^{V}$, $W^{O}$) with $r=128$, $\alpha=256$, and additionally tune scaling-related modules for SALAD (see Fig. 3). Training is performed on 4 GPUs with total batch size 8, AdamW optimizer, and learning rate $1\mathrm{e}{-4}$.
Wan2.1-1.3B (480p) Wan2.1-1.3B (720p) Wan2.1-14B (480p)
Model Wan2.1-1.3B Wan2.1-1.3B Wan2.1-14B
Resolution $480\times 832$ $720\times 1280$ $480\times 832$
Approx. Tokens $\sim$30k $\sim$100k $\sim$30k
Tuning Steps 1,600 1,000 600
GPU Hours (est.) $\sim$20 $\sim$30 $\sim$20
Optimizer AdamW AdamW AdamW
Learning rate $1\mathrm{e}{-4}$ $1\mathrm{e}{-4}$ $1\mathrm{e}{-4}$
LoRA rank $r$ / $\alpha$ 128 / 256 128 / 256 128 / 256
Batch size 8 (4 GPUs) 8 (4 GPUs) 8 (4 GPUs)
Training samples 2,000 (Mixkit) 2,000 (Mixkit) 2,000 (Mixkit)
0.A.7.2 Training Cost
SALAD requires only 2,000 training videos and a relatively small number of optimization steps. For Wan2.1-1.3B at 480p, the entire tuning process takes approximately 20 GPU hours with a batch size of 8. Even at higher resolution (720p), the tuning cost remains modest at around 30 GPU hours. For the larger Wan2.1-14B model at 480p, SALAD requires only 600 training steps (about 20 GPU hours).
Overall, SALAD enables efficient adaptation of sparse-linear hybrid attention in Video Diffusion Transformers with a substantially lower computational cost than approaches that pretrain sparse-attention architectures from scratch.
For reference, VSA [zhang2025vsa] is pretrained on 80,000 videos using 32 NVIDIA H100 GPUs, while VMoBA [wu2025vmoba] leverages the Koala-36M dataset (36 million video clips) and 104 NVIDIA H800 GPUs under the token sequence length of 33K. SLA [zhang2025sla], a recent linear-attention-based approach for high-sparsity adaptation, is tuned on 20,000 self-collected videos—ten times more than SALAD. SLA also requires a batch size of 64 for tuning. Under our experimental setting, SLA fails to recover the performance of the original dense model as reported in Tab. 1, whereas SALAD achieves comparable or superior results at a fraction of the resource cost. These findings underscore SALAD’s tuning efficiency for sparse adaptation in large-scale video diffusion models.
0.A.8 More Qualitative Results
@@
To further assess the visual quality and temporal consistency achieved by SALAD, we present additional qualitative comparisons against several representative baselines in this section. We evaluate SALAD with other baselines on Wan 2.1-1.3B. As illustrated in Figure 14 and Figure 15, SALAD consistently preserves high-fidelity appearance details while generating temporally coherent motion, often matching or surpassing the visual quality produced by dense-attention counterparts.